Solve the system of linear equation using matrices
x+y=1
x-7y=17

I wanted to double chek it but i got (3,-2) is it right? if not please correct me.

Answer by Alan3354(69443)

If you sub for x and y, you'll find that it's correct.
---------
x+y=1
3 - 2 = 1
------------
x-7y=17
3 - 7(-2) = 17